### v1

@@ 2 System model
We consider a network that consists of multi-antenna APs and macro BSs, and single-antenna MSs. The APs are connected by means of fronthaul links to a CPU where data encoding/decoding is performed. Depending on the type of envisaged cooperation between the BSs and the APs, also the BSs can be connected through fronthaul links to a CPU. All communications take place on the same frequency band, i.e. uplink and downlink are separated through time-division-duplex (TDD), and there is perfect synchronization in the system.
While the MSs are equipped with a single antenna, each AP is equipped with an array of $N_{\rm AP}$ antennas. We assume that BS sites are actually made of three sectors and the sectors are treated as independent base stations, i.e., we have $L$ BSs arranged in $L/3$ BS sites. Each sector is equipped with a uniform linear array (ULA) made of $N_{\rm BS}$ antennas with half wavelength spacing.
@@ 2.1 Network deployment scenarios
Although the subsequent derivations hold for any arbitrary configuration for the deployment of BSs and of the APs, our analysis will refer to a network deployment where the three-sector macro BS sites are placed according to the classical hexagonal layout, while APs are deployed according to two possible strategies, namely either uniform deployment or cell-edge deployment. The former case accounts for the situation where we want to densify the network by bringing network terminating points closer to the users so as to have the capability to further push the system capacity; the latter deployment, instead, aims similarly at increasing the network throughput, but exploiting APs to provide better coverage at the critical cell-edge, so as to increase the system fairness across users. Figure 1 shows an instance of the two considered deployments. In particular, in the figure we also highlight users that are at a distance from a BS that is smaller than one third of the inter-site distance (ISD) and the remaining users. The former are called cell-inside users, while the latter will be termed cell-edge user. Such distinction is made to study the effects on the two sets of the network deployment schemes considered in the following.
@@
    A deployment with $L$ MC-mMIMO BSs, arranged over equally spaced sites with 120 deg sectors, and with no-BS cooperation. This scenario will be indicated as ”MC-mMIMO”.
@@
    An heterogeneous non cooperative network where a MC-mMIMO and a CF-mMIMO system co-exist; it is assumed here that each UE can connect either to one AP or to one BS, and that there is no cooperation. This scenario resembles a network with base stations of two different kinds and will be named ”HET-noCoop”.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with horizontal cooperation; it is here assumed that each UE can be cooperatively served by either a certain number of BSs or by a certain number of APs. However, there is no joint cooperation among BSs and APs. This scenario will be indicated as ”HORIZONTAL-Coop”.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with full cooperation; it is here assumed that all the UEs are cooperatively and jointly served by several APs and several BSs. This is a fully-cooperative scenario and will be labeled as ”FULL-Coop”.
@@ 2.2 Channel Model
We consider a typical channel model for sub-6GHz carrier frequencies, based on the rich scattering environment assumption. We denote by $\mathbf{g}_{k,m}\in\mathbb{C}^{N_{\rm AP}}$ ($\mathbf{h}_{k,\ell}\in\mathbb{C}^{N_{\rm BS}}$) the channel between the $k$-th user and the $m$-th AP ($\ell$-th BS). We assume that
$\mathbf{g}_{k,m}=\sqrt{\beta_{k,m}}\overline{\mathbf{g}}_{k,m},\quad\mathbf{h}_{k,\ell}=\sqrt{\rho_{k,\ell}}\overline{\mathbf{h}}_{k,\ell}.$ (1)
In the above equation, $\beta_{k,m}$ and $\rho_{k,\ell}$ are scalar coefficients modeling the channel path loss and shadowing effects between the $k$-th UE and the $m$-th AP andthe $\ell$-th BS, respectively. Moreover $\overline{\mathbf{g}}_{k,m}$ and $\overline{\mathbf{h}}_{k,\ell}$ are an $N_{\rm AP}$-dimensional vector and an $N_{\rm BS}$-dimensional vector, respectively, whose entries are i.i.d ${\cal CN}(0,1)$ random variates (RVs) modeling the fast fading. Further details on the channel model will be given in Section VII on the numerical results.
@@ 2.3 Uplink Training
The dimension in time/frequency samples of the channel coherence length is denoted by $\tau_{c}$, and the dimension of the uplink training phase is expressed as $\tau_{p}<\tau_{c}$. We assume that the MSs transmit their pilot sequences and the channel estimation is performed locally at the APs and at the BSs in the system. The pilot sequences transmitted by the users, say $\bm{\phi}_{1},\ldots,\bm{\phi}_{K}$, are chosen in a set of $\tau_{p}$ orthogonal sequences $\mathcal{P}_{\tau_{p}}=\left\{\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{\tau_{p}}\right\}$, where $\bm{\psi}_{i}$ is the $i$-th $\tau_{p}$-dimensional column sequence and $\|\bm{\psi}_{i}\|^{2}=1$, $\forall\,i=1,\ldots,\tau_{p}$. In Section 2.4 we discuss the pilot assignment algorithm used in this paper.
The $m$-th AP, during the training phase, receives the signal $\mathbf{Y}_{m}^{(a)}\in\mathbb{C}^{N_{\rm AP}\times\tau_{p}}$, and, assuming knowledge of the large-scale fading coefficients $\beta_{k,m}$, estimates the fast fading component of the channel from the $k$-th user using the linear minimum-mean-square-error (LMMSE) criterion [26] as follows:
$\widehat{\mathbf{g}}_{k,m}=\frac{\sqrt{\eta_{k}}\beta_{k,m}}{\displaystyle\sum_{i=1}^{K}\eta_{i}\beta_{i,m}\left|\bm{\phi}_{i}^{H}\bm{\phi}_{k}\right|^{2}+\sigma^{2}_{w}}\widehat{\mathbf{y}}_{k,m}^{(a)}=\alpha_{k,m}\widehat{\mathbf{y}}_{k,m}^{(a)}\;,$ (2)
with $\widehat{\mathbf{y}}_{k,m}^{(a)}=\mathbf{Y}_{m}^{(a)}\bm{\phi}_{k}$. The channel estimation error is given by $\widetilde{\mathbf{g}}_{k,m}=\mathbf{g}_{k,m}-\widehat{\mathbf{g}}_{k,m}$, and the estimate and the estimation error are independent [26]. They are distributed as $\widehat{\mathbf{g}}_{k,m}\sim\mathcal{CN}\left(\mathbf{0}_{N_{\rm AP}},\gamma_{k,m}\mathbf{I}_{N_{\rm AP}}\right)$, and $\widetilde{\mathbf{g}}_{k,m}\sim\mathcal{CN}\left(\mathbf{0}_{N_{\rm AP}},(\beta_{k,m}-\gamma_{k,m})\mathbf{I}_{N_{\rm AP}}\right)$, respectively, where $\gamma_{k,m}$ is the mean-square of the estimate, i.e., for any antenna element $n=1,\ldots,N_{\rm AP}$,
$\gamma_{k,m}=\mathbb{E}\left[\left|\left(\widehat{\mathbf{g}}_{k,m}\right)_{n}\right|^{2}\right]=\sqrt{\eta_{k}}\beta_{k,m}\alpha_{k,m}\,.$ (3)
@@
where $\mathbf{W}_{\ell}^{(b)}\in\mathbb{C}^{N_{\rm BS}\times\tau_{p}}$ contains the thermal noise contribution and out-of-cell interference at the $\ell$-th BS, with i.i.d. ${\cal CN}(0,\sigma^{2}_{w})$ RVs as entries. From the observable $\mathbf{Y}_{\ell}^{(b)}$, the $\ell$-th BS can estimate the channel vectors $\left\{\mathbf{h}_{k,\ell}\right\}_{k=1}^{K}$ based on the statistics $\widehat{\mathbf{y}}_{k,\ell}^{(b)}=\mathbf{Y}_{\ell}^{(b)}\bm{\phi}_{k}$ through the LMMSE approach. Assuming again knowledge of the coefficients $\rho_{k,\ell}$, the LMMSE channel estimate of $\mathbf{h}_{k,\ell}$ can be written as [26]
$\widehat{\mathbf{h}}_{k,\ell}=\frac{\sqrt{\eta_{k}}\rho_{k,\ell}}{\displaystyle\sum_{i=1}^{K}\eta_{i}\rho_{i,\ell}\left|\bm{\phi}_{i}^{H}\bm{\phi}_{k}\right|^{2}+\sigma^{2}_{w}}\widehat{\mathbf{y}}_{k,\ell}^{(b)}=\xi_{k,\ell}\widehat{\mathbf{y}}_{k,\ell}^{(b)}\;.$ (5)
The channel estimation error is given by $\widetilde{\mathbf{h}}_{k,\ell}=\mathbf{h}_{k,\ell}-\widehat{\mathbf{h}}_{k,\ell}$, and the estimate and the estimation error are independent [26]. They are distributed as $\widehat{\mathbf{h}}_{k,\ell}\sim\mathcal{CN}\left(\mathbf{0}_{N_{\rm BS}},\delta_{k,\ell}\mathbf{I}_{N_{\rm BS}}\right)$, and $\widetilde{\mathbf{h}}_{k,\ell}\sim\mathcal{CN}\left(\mathbf{0}_{N_{\rm BS}},(\rho_{k,\ell}-\delta_{k,\ell})\mathbf{I}_{N_{\rm BS}}\right)$, respectively, where $\delta_{k,\ell}$ is the mean-square of the estimate, i.e., for any antenna element $n=1,\ldots,N_{\rm BS}$,
$\delta_{k,\ell}=\mathbb{E}\left[\left|\left(\widehat{\mathbf{h}}_{k,\ell}\right)_{n}\right|^{2}\right]=\sqrt{\eta_{k}}\rho_{k,\ell}\xi_{k,\ell}\,.$ (6)
@@ 2.4 Pilot Assignment Algorithm
We now discuss how the $\tau_{p}$ orthogonal pilot sequences in $\mathcal{P}_{\tau_{p}}$ are to be shared among the $K$ users, with $K\gg\tau_{p}$. In general, the distance between the users using the same pilot sequence should be as large as possible in order to reduce the impact of pilot contamination on the system performance. It is thus assumed that the users’ positions are known to the network operator or it can be estimated. The proposed procedure uses the well-known $k$-means clustering method [27], i.e., an iterative algorithm that is able to partition users into disjoint clusters. Defining the centroid of each cluster as the mean of the positions of the APs in the cluster, the algorithm, accepting as input the users positions and the number of pilot sequences is summarized in Algorithm 1 and operates as follows:
@@
    $\lceil K/\tau_{p}\rceil$ centroids are chosen so that they approximately form a regular grid over the considered area, i.e., the so-called parameter ”$k$” in the $k$-means clustering is $\lceil K/\tau_{p}\rceil$.
@@
    Repeat steps [b.] and [c] until the positions of the centroids converge.
@@
Regarding step [e.], its aim is to ensure that users that are assigned the same pilot sequence are not too close. The methodology that we propose here, based on the latitude only of the users and not on their full 2D coordinates, is clearly suboptimal, but has been tested to represent a good trade-off between complexity and performance.

### v2

@@ 2 System model
We consider a network that consists of multi-antenna APs and macro-BSs, and single-antenna MSs. The APs are connected by means of fronthaul links to a CPU where data encoding/decoding is performed. Depending on the type of cooperation envisaged between the BSs and the APs, the BSs can also be connected through fronthaul links to a CPU. All communications take place on the same frequency band, i.e. uplink and downlink are separated through time-division-duplex (TDD), and there is perfect synchronization in the system.
Although the MSs are equipped with a single antenna, each AP is equipped with a uniform linear array (ULA) made of $N_{\rm AP}$ antennas with half-wavelength spacing. We assume that BS sites are actually made of three sectors and the sectors are treated as independent base stations, i.e., we have $L$ BSs arranged in $L/3$ BS sites. Each sector is equipped with a ULA made of $N_{\rm BS}$ antennas with half-wavelength spacing.
In Table 1 we report the definition of the main symbols used in the document.
Table 1: Definition of the main symbols used in the paper
Name Meaning
$L$ number of BSs (sectors)
$M$ number of APs
$K$ number of users
$N_{\rm BS}$ number of antennas at the BSs
$N_{\rm AP}$ number of antennas at the APs
$\mathbf{h}_{k,\ell}$ channel between the $k$-th user and the $\ell$-th BS
$\mathbf{g}_{k,m}$ channel between the $k$-th user and the $m$-th AP
$\rho_{k,\ell}$ scalar coefficients modeling the channel path loss and shadowing effects between the $k$-th UE and the $\ell$-th BS
$\beta_{k,m}$ scalar coefficients modeling the channel path loss and shadowing effects between the $k$-th UE and the $m$-th AP
$\tau_{p}$ length of the training phase in samples
$\bm{\phi}_{k}$ pilot sequence transmitted by the $k$-th user
$\eta_{k}$ power transmitted by the $k$-th user during the training phase
$\widehat{\mathbf{h}}_{k,\ell}$ estimate of the channel between the $k$-th user and the $\ell$-th BS
$\widehat{\mathbf{g}}_{k,m}$ estimate of the channel between the $k$-th user and the $m$-th AP
$b_{k,\ell}$ the binary variates which is 1 if the $k$-th user is served by the $\ell$-th BS and 0 otherwise
$a_{k,m}$ the binary variates which is 1 if the $k$-th user is served by the $m$-th AP and 0 otherwise
$\mathbf{w}_{k,\ell}^{(b)}$ unit-norm beamforming vector used at the $\ell$-th BS for the downlink transmission to the $k$-th user
$\mathbf{w}_{k,m}^{(a)}$ unit-norm beamforming vector used at the $m$-th AP for the downlink transmission to the $k$-th user
$\eta_{k,\ell}^{(b)}$ positive scalar coefficient controlling the power transmitted by the $\ell$-th BS to the $k$-th user
$\eta_{k,m}^{(a)}$ positive scalar coefficient controlling the power transmitted by the $m$-th AP to the $k$-th user
$P_{\ell,{\rm max}}^{(b)}$ maximum power available at the $\ell$-th BS
$P_{m,{\rm max}}^{(a)}$ maximum power available at the $m$-th AP
@@ 2.1 Network deployment scenarios
Although the subsequent derivations hold for any arbitrary configuration for the deployment of BSs and of the APs, our analysis will refer to a network deployment in which the three-sector macro BS sites are placed according to the classical hexagonal layout, while APs are deployed according to two possible strategies, namely uniform deployment or cell-edge deployment. The former case accounts for the situation where we want to densify the network by bringing network terminating points closer to the users so as to have the capability to further push the system capacity; the latter deployment, instead, aims similarly at increasing the network throughput, but exploiting APs to provide better coverage at the critical cell-edge, so as to increase the system fairness across users. Figure 1 shows an example of the two deployments considered. In particular, in the figure we also highlight users that are at a distance from a BS that is smaller than one third of the inter-site distance (ISD) and the remaining users. The former are called cell-inside users, while the latter will be termed cell-edge users. This distinction is made to study the effects on the two sets of the network deployment schemes considered below.
@@
    A deployment with $L$ MC-mMIMO BSs, arranged over equally spaced sites with 120 deg sectors, and with no-BS cooperation. This scenario will be indicated as “MC-mMIMO”.
@@
    An heterogeneous non-cooperative network where a MC-mMIMO and a CF-mMIMO system co-exist; it is assumed here that each UE can connect either to one AP or to one BS, and that there is no cooperation. This scenario resembles a network with base stations of two different kinds and will be named “HET-noCoop”.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with horizontal cooperation; it is assumed here that each UE can be cooperatively served by either a certain number of BSs or by a certain number of APs. However, there is no joint cooperation among BSs and APs. This scenario will be indicated as “HORIZONTAL-Coop”.
@@
    A MC-mMIMO network co-existing with a CF-mMIMO network with full cooperation; it is assumed here that all the UEs are cooperatively and jointly served by several APs and several BSs. This is a fully cooperative scenario and will be labeled as “FULL-Coop”.
@@ 2.2 Channel Model
The channel model under consideration in the frequency domain includes a combination of a large-scale fading coefficient (covering shadowing and path-loss effects) and a small-scale fading vector. We generally assume that the latter follows a Ricean distribution to represent the potential existence of a line-of-sight (LOS) component over short-distance links. [14]. More precisely, $\mathbf{g}_{k,m}\in\mathbb{C}^{N_{\rm AP}}$ ($\mathbf{h}_{k,\ell}\in\mathbb{C}^{N_{\rm BS}}$) denotes the channel between the $k$-th user and the $m$-th AP ($\ell$-th BS), and is expressed as
$\mathbf{g}_{k,m}=\sqrt{\frac{\beta_{k,m}}{K_{k,m}^{(a)}+1}}\left[\sqrt{K_{k,m}^{(a)}}e^{j\vartheta_{k,m}}\mathbf{a}_{AP}\left(\theta_{k,m}\right)+\overline{\mathbf{g}}_{k,m}\right]\;,$ (1)
$\mathbf{h}_{k,\ell}=\sqrt{\frac{\rho_{k,\ell}}{K_{k,\ell}^{(b)}+1}}\left[\sqrt{K_{k,\ell}^{(b)}}e^{j\varphi_{k,\ell}}\mathbf{a}_{BS}\left(\phi_{k,\ell}\right)+\overline{\mathbf{h}}_{k,\ell}\right]\;.$ (2)
In the above equations, $\beta_{k,m}$ ($\rho_{k,\ell}$) is a scalar coefficient modeling the large-scale fading, i.e. the channel path loss and shadowing effects, between the $k$-th UE and the $m$-th AP ($\ell$-th BS). Moreover, $\overline{\mathbf{g}}_{k,m}$ ($\overline{\mathbf{h}}_{k,\ell}$) is an $N_{\rm AP}$-dimensional vector ($N_{\rm BS}$-dimensional vector), whose entries are i.i.d. ${\cal CN}(0,1)$ random variates (RVs) modeling the fast fading. The parameter $K_{k,m}^{(a)}$ $\left(K_{k,\ell}^{(b)}\right)$ is the Rician $K$-factor for the channel from the $k$-th user to the $m$-th AP ($\ell$-th BS); this factor usually depends on the probability that there is a LOS link between the user and the AP (BS). Letting $p_{\rm LOS}^{(a)}\left(d_{k,m}^{(a)}\right)$ $\left(p_{\rm LOS}^{(b)}\left(d_{k,\ell}^{(b)}\right)\right)$ denote the probability that there is a LOS link between the $k$-th user and the $m$-th AP ($\ell$-th BS), the Ricean factors are evaluated as [14],
$K_{k,m}^{(a)}=\frac{p_{\rm LOS}^{(a)}\left(d_{k,m}^{(a)}\right)}{1-p_{\rm LOS}^{(a)}\left(d_{k,m}^{(a)}\right)}\,,$ (3)
and
$K_{k,\ell}^{(b)}=\frac{p_{\rm LOS}^{(b)}\left(d_{k,\ell}^{(b)}\right)}{1-p_{\rm LOS}^{(b)}\left(d_{k,\ell}^{(b)}\right)}\,,$ (4)
respectively. The quantities $\vartheta_{k,m}$ and $\varphi_{k,\ell}$ in (1) and (2) follow a uniform distribution in $[0,2\pi]$, and denote the random phase offsets associated to the direct path from the $k$-th user to the $m$-th AP and to the $\ell$-th BS, respectively. Finally, $\mathbf{a}_{AP}\left(\theta_{k,m}\right)\in\mathbb{C}^{N_{\rm AP}}$ and $\mathbf{a}_{BS}\left(\phi_{k,\ell}\right)\in\mathbb{C}^{N_{\rm BS}}$ represent the steering vectors characterizing the direct path between the $k$-th user and the $m$-th AP or the $\ell$-th BS, respectively, and follow the model referenced in  [14]. More details on the channel model will be given in Section 7.7.1 and in the Appendix.
@@ 2.3 Uplink Training
The dimension in time/frequency samples of the channel coherence length is denoted by $\tau_{c}$, and the dimension of the uplink training phase is expressed as $\tau_{p}<\tau_{c}$. We assume that the MSs transmit their assigned pilot sequence and that the channel estimation is performed locally at the APs and at the BSs in the system. The pilot sequences transmitted by the users, say $\bm{\phi}_{1},\ldots,\bm{\phi}_{K}$, are chosen in a set of $\tau_{p}$ orthogonal sequences $\mathcal{P}_{\tau_{p}}=\left\{\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{\tau_{p}}\right\}$, where $\bm{\psi}_{i}$ is the $i$-th $\tau_{p}$-dimensional column sequence and $\|\bm{\psi}_{i}\|^{2}=1$, $\forall\,i=1,\ldots,\tau_{p}$. In Section 2.2.4 we discuss the pilot assignment algorithm used in this paper.
The $m$-th AP, during the training phase, receives the signal $\mathbf{Y}_{m}^{(a)}\in\mathbb{C}^{N_{\rm AP}\times\tau_{p}}$,
$\mathbf{Y}_{m}^{(a)}=\displaystyle\sum_{i=1}^{K}\displaystyle\sqrt{\eta_{i}}\mathbf{g}_{i,m}\bm{\phi}_{i}^{H}+\mathbf{W}_{m}^{(a)}\;,$ (5)
where $\eta_{i}=\tau_{p}\overline{\eta}_{i}$ is the power transmitted by the $i$-th user during the uplink training phase, and $\mathbf{W}_{m}^{(a)}\in\mathbb{C}^{N_{\rm AP}\times\tau_{p}}$ contains the thermal noise contribution and out-of-cell interference at the $\ell$-th BS, with i.i.d. ${\cal CN}(0,\sigma^{2}_{w})$ RVs as entries. Assuming knowledge of the large-scale fading coefficients $\beta_{k,m}$, of the vectors $\mathbf{a}_{AP}\left(\theta_{k,m}\right)$, and of the Rician factors $K_{k,m}^{(a)}\,\forall\;m,k$, the estimates of the of the channel ${\mathbf{g}}_{k,m}$ by using the LMMSE criterion [30] can be obtained as follows:
$\widehat{\mathbf{g}}_{k,m}=\mathbf{D}_{k,m}^{(a)}\widehat{\mathbf{y}}_{k,m}^{(a)}\;,$ (6)
with $\widehat{\mathbf{y}}_{k,m}^{(a)}=\mathbf{Y}_{m}^{(a)}\bm{\phi}_{k}$ and
$\begin{array}[]{lllll}\mathbf{D}_{k,m}^{(a)}=\sqrt{\eta_{k}}\mathbf{G}_{k,m}\mathbf{B}_{k,m}^{-1}\in\mathbb{C}^{N_{\rm AP}\times N_{\rm AP}}\,,\\ \mathbf{G}_{k,m}=\frac{\beta_{k,m}}{K_{k,m}^{(a)}+1}\left[K_{k,m}^{(a)}\mathbf{a}_{AP}\left(\theta_{k,m}\right)\mathbf{a}_{AP}^{H}\left(\theta_{k,m}\right)+\mathbf{I}_{N_{\rm AP}}\right]\,,\\ \mathbf{B}_{k,m}=\sum_{i\in\mathcal{K}}{\eta_{i}\mathbf{G}_{i,m}\left|\bm{\phi}_{i}^{H}\bm{\phi}_{k}\right|^{2}}+\sigma^{2}_{w}\mathbf{I}_{N_{\rm AP}}.\end{array}$
@@
where $\mathbf{W}_{\ell}^{(b)}\in\mathbb{C}^{N_{\rm BS}\times\tau_{p}}$ contains the thermal noise contribution and out-of-cell interference at the $\ell$-th BS, with i.i.d. ${\cal CN}(0,\sigma^{2}_{w})$ RVs as entries. From the observable $\mathbf{Y}_{\ell}^{(b)}$, the $\ell$-th BS can estimate the channel vectors $\left\{\mathbf{h}_{k,\ell}\right\}_{k=1}^{K}$ based on the statistics $\widehat{\mathbf{y}}_{k,\ell}^{(b)}=\mathbf{Y}_{\ell}^{(b)}\bm{\phi}_{k}$ through the LMMSE approach. Assuming again knowledge of the large-scale fading coefficients $\rho_{k,\ell}$, of the vectors $\mathbf{a}_{BS}\left(\phi_{k,\ell}\right)$, and of the Rician factors $K_{k,\ell}^{(b)}\,\forall\;\ell,k$, the estimates of the channel ${\mathbf{h}}_{k,\ell}$ by using the LMMSE criterion [30] can be obtained as follows:
$\widehat{\mathbf{h}}_{k,\ell}=\mathbf{D}_{k,\ell}^{(b)}\widehat{\mathbf{y}}_{k,\ell}^{(b)}\;,$ (8)
with
$\begin{array}[]{lllll}\mathbf{D}_{k,\ell}^{(b)}=\sqrt{\eta_{k}}\mathbf{H}_{k,\ell}\mathbf{C}_{k,\ell}^{-1}\in\mathbb{C}^{N_{\rm BS}\times N_{\rm BS}}\,,\\ \mathbf{H}_{k,\ell}=\frac{\rho_{k,\ell}}{K_{k,\ell}^{(b)}+1}\left[K_{k,\ell}^{(b)}\mathbf{a}_{BS}\left(\phi_{k,\ell}\right)\mathbf{a}_{BS}^{H}\left(\phi_{k,\ell}\right)+\mathbf{I}_{N_{\rm BS}}\right]\,,\\ \mathbf{C}_{k,\ell}=\sum_{i\in\mathcal{K}}{\eta_{i}\mathbf{H}_{i,\ell}\left|\bm{\phi}_{i}^{H}\bm{\phi}_{k}\right|^{2}}+\sigma^{2}_{w}\mathbf{I}_{N_{\rm BS}}.\end{array}$
@@ 2.4 Pilot Assignment Algorithm
We now discuss how the $\tau_{p}$ orthogonal pilot sequences in $\mathcal{P}_{\tau_{p}}$ are to be shared among the $K$ users, with $K\gg\tau_{p}$. In general, the distance between users using the same pilot sequence should be as large as possible to reduce the impact of pilot contamination on the system performance. It is thus assumed that the users’ positions are known to the network operator, or they can be estimated. The proposed procedure uses the well-known $k$-means clustering method [31], i.e., an iterative algorithm that is able to partition users into disjoint clusters. Defining the centroid of each cluster as the mean of the positions of the APs in the cluster, the algorithm, accepting as input the users positions and the number of pilot sequences, is summarized in Algorithm 1 and operates as follows:
@@
    $\lceil K/\tau_{p}\rceil$ centroids are chosen so that they approximately form a regular grid over the considered area, i.e., the so-called parameter “$k$” in the $k$-means clustering is $\lceil K/\tau_{p}\rceil$.
@@
    Repeat steps [b.] and [c.] until the positions of the centroids converge.
@@
Regarding step [e.], its aim is to ensure that users that are assigned the same pilot sequence are not too close. The methodology that we propose here, based only on the latitude of the users and not on their full 2D coordinates, is clearly suboptimal, but has been tested to represent a good trade-off between complexity and performance.